When a sigle slit is replaced by the double slit arrangement, the intensity of central maximum is increased by?
1
Expert's answer
2020-06-08T10:30:17-0400

Since,

"I_{double \:slit}=I_o\cos^2(\frac{\Delta \phi}{2})"

Where,

"\Delta \phi=\frac{2\pi}{\lambda}d\sin(\theta)"

Now, Central maximum occurs when "d\sin(\theta)=\lambda" ,thus from above formula we get,

"I_{double \:slit}=I_o\cos^2(\frac{2\pi}{2})=I_o\hspace{1cm}(1)"

And

"I_{single \: slit}=I_o\bigg(\frac{\sin(\beta)}{\beta}\bigg)^2"

where,

"\beta=\frac{\pi}{\lambda}a\sin(\theta)"

Thus, we have central maximum when,"\theta\rightarrow 0\implies \beta\rightarrow0" ,thus"\frac{\sin(\beta)}{\beta}\rightarrow 1" ,

Hence,

"I_{single \: slit}=I_o\hspace{1cm}(2)"

Therefore, from equations (1) and (2) we get,

Intensity for both the slit at central maximum is equal.